OpenStudy (anonymous):

A tank can be filled by a pipe in 7 hours. While a drain pipe can empty the tank in 11 hours. If the first 4 hours, the drain was left open while the pipe is filling it. How long will the tank be filled with water?

OpenStudy (hba):

Let x = time (hours) for second pipe to fill alone then 3(1/4 + 1/x) = 1 multiplying both sides by 4x: 3(x + 4) = 4x 3x + 12 = 4x 12 hours = x

OpenStudy (hba):

Another solution: 1st pipe rate = 1/4 job/hr --- Together rate = 1/3 job/hr ---- 2nd pipe rate = 1/x job/hr ---------------------------------- Equation: rate + rate = together rate 1/4 + 1/x = 1/3 Multiply thru by 12x to get: 3x + 12 = 4x x = 12 hrs (time required for the 2nd pipe to do the job alone) @stupidinmath